(1/x-1)-(1/x+1)=(2/x^2-1)
list the value of restriction on the variable and solve the equation

Answer by oscargut(2103)

(1/x-1)-(1/x+1)=(2/x^2-1)
Restrictions:
x canīt be 1 or -1
(1/x-1)-(1/x+1)=(2/x^2-1)
[(x+1)-(x-1)]/(x^2-1) = 2/(x^2-1)
(x+1)-(x-1)= 2
2 = 2
Solution set = {x/ x is real} - {-1,1}
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com